Calculate the change in entropy when 12g of water at -10oC is raised to 100oC and convert to steam at 100oC. Assume that specific heat of water cwater = 4190 J·g-1·K-1, heat of vaporization Lv = 2256 kJ/kg,specific heat of ice cice = 2220 J·g-1·K-1, heat of fusion LF = 334 kJ/kg.
1
Expert's answer
2019-06-10T09:41:03-0400

Using SI:

"T_1=-10^oC+273=263 K" - tempersture of starting heating

"T_2=0^oC+273=273 K" - temperature of ice melting

"T_3=100^oC+273=373 K" - temperature of converting to steam

"m=0.012 kg" - mass of water

"c_{water}= 4190 Jkg^{-1}K^{-1}" - specific heat of water //mistake in statment "g \to kg"

"c_{ice}= 2220 Jkg^{-1}K^{-1}"- specific heat of ice //mistake in statment "g \to kg"

"L_v=2256*10^3 Jkg^{-1}" - heat of vaporization

"L_f=334*10^3 Jkg^{-1}" - heat of fusion


For specific heat: "C=c*m"

For phase transition: "\Delta H=L*m"


The process could be divided into 4 parts:

Heatind of ice to 0 degrees Celsius. Change of entropy:

"\Delta S_1=Cln(T_2/T_1)=c_{ice}mln(T_2/T_1)"


Melting of the ice at constant temperature. Chenge of entropy:


"\Delta S_2=\Delta H_f/T_1=L_fm/T_1"

Heatind of water from 0 to 100 degrees Celsius. Change of entropy:


"\Delta S_3=Cln(T_3/T_2)=c_{water}mln(T_3/T_2)"

Vaporization. Change of entropy:


"\Delta S_4=\Delta H_v/T_3=L_vm/T_3"

Total change of entropy:


"\Delta S=\Delta S_1+\Delta S_2+\Delta S_3+\Delta S_4 \implies"

"\Delta S=c_{ice}mln(T_2/T_1)+L_fm/T_1+c_{water}mln(T_3/T_2)+L_vm/T_3"

Using numbers:


"\Delta S=103.96 J/K"

Answer: "\Delta S=103.96 J/K"